Which of the following listed below a subgroup of G, if H and K are two subgroups of a group G?
a.H\\(\\mathrm{\\cap}\\)K
b.H/K
c.H-K
d.H\\(\\cup\\)K
1
Expert's answer
2020-06-02T18:49:49-0400

Correct option is (a).

Reason:


Let fix notation: "\\\leq" denote subgroup.

Now, we have given "G" is a group,"H\leq G \&K\leq G" .

Claim:

"H\cap K\leq G"

Proof:

It is sufficient to show "H\cap K \leq G" if we show "H\cap K" closed under multiplication and taking inverse in it.

Let, for every "h,k\in H\cap K \implies h\in H \& h\in K" also "k\in H \& k\in K" ,thus


"hk\in H\&hk\in K \hspace{1cm}(\because H,K\leq G)"

Hence,

"hk\in H\cap K"

Which implies "H\cap K" is closed under multiplication.

Now, consider

"\forall h\in H\cap K\implies h\in H\&h\in K"

Since, "H,K\leq G" ,

"h^{-1}\in H\&h^{-1}\in K\implies h^{-1}\in H\cap K"

Thus, "H\cap K" is closed under taking inverse.

Therefore,

"H\cap K \leq G"

For remaining 3 options we observe that if we choose "G=S_3" and "H=\{e,(12)\}\&K=\{e,(13)\}" , does not qualify the reaming options to be subgroup of "G" .